### v1

# Federated Retrieval-Augmented Generation: A Systematic Mapping Study

Abstract
Federated Retrieval-Augmented Generation (Federated RAG) combines Federated Learning (FL),which enables distributed model training without exposing raw data, with Retrieval-Augmented Generation (RAG), which improves the factual accuracy of language models by grounding outputs in external knowledge. As large language models are increasingly deployed in privacy-sensitive domains such as healthcare, finance, and personalized assistance, Federated RAG offers a promising framework for secure, knowledge-intensive natural language processing (NLP). To the best of our knowledge, this paper presents the first systematic mapping study of Federated RAG, covering literature published between 2020 and 2025. Following Kitchenham’s guidelines for evidence-based software engineering, we develop a structured classification of research focuses, contribution types, and application domains. We analyze architectural patterns, temporal trends, and key challenges, including privacy-preserving retrieval, cross-client heterogeneity, and evaluation limitations. Our findings synthesize a rapidly evolving body of research, identify recurring design patterns, and surface open questions, providing a foundation for future work at the intersection of RAG and federated systems.
[…]
intermediate/.style=font=,
Federated Retrieval-Augmented Generation: A Systematic Mapping Study
Abhijit Chakrabortythanks: Corresponding author1  and Chahana Dahal2  and Vivek Gupta11footnotemark: 11 1Arizona State University   2Westminster University {achakr40,vgupt140}@asu.edu
1 Introduction
Large language models (LLMs) are increasingly deployed in domains such as healthcare, finance, and personalized assistance, raising concerns about data privacy, ownership, and regulatory compliance. In response, Federated Learning (FL) has emerged as a compelling paradigm for training models across distributed clients without exchanging raw data (Kairouz et al., 2021). In parallel, Retrieval-Augmented Generation (RAG) (Lewis et al., 2020a) enhances LLMs by grounding their outputs in dynamically retrieved external knowledge, reducing hallucinations and improving the accuracy of facts (Lewis et al., 2020b).
FL enables distributed model training across devices or institutions. Each client trains a model locally and only shares model updates (e.g., gradients), not raw data with each other. Common techniques like Federated Averaging (FedAvg) (McMahan et al., 2017) and secure aggregation (Bonawitz et al., 2017) help combine these updates safely. FL has been applied in healthcare, edge devices (e.g., smartphones), and finance, where data privacy is critical (Kairouz et al., 2021).
RAG is a hybrid method that combines a text retriever with a generator. Before producing a response, the model first retrieves relevant text from an external knowledge base or document store. This process helps reduce hallucinations and improves factual consistency in the generated output. Popular implementations include the original RAG model (Lewis et al., 2020b) and Fusion-in-Decoder (FiD) (Izacard and Grave, 2020), which demonstrate how integrating retrieval can enhance open-domain question answering.
At the intersection of these paradigms lies Federated RAG, a hybrid approach that enables LLMs to access distributed knowledge sources in a privacy-preserving manner. It combines the strengths of data localization, personalized retrieval, and context-aware generation. Conceptually, Federated RAG builds on earlier federated search methods (Shokouhi and Si, 2011), which aggregated results from siloed sources without centralized indexing, and extends them to support complex generative tasks.
As LLMs have grown in capability, driven by innovations such as the Transformer architecture (Vaswani et al., 2017), pre-training methods (Devlin et al., 2019; Radford et al., 2019; Brown et al., 2020), and prompting strategies (Schick and Schütze, 2021; Liu et al., 2021; Debnath et al., 2025), so has the demand for integrating them with heterogeneous private data contexts. Since 2019, the parallel evolution of LLMs, RAG, and FL has laid the groundwork for Federated RAG to emerge as both a viable and increasingly necessary framework.
In this paper, we present the first systematic mapping study of Federated RAG, following the methodology of evidence-based software engineering proposed by Kitchenham (Kitchenham et al., 2011). Our objective is to map how retrieval-augmented generation is being adapted and deployed across federated architectures, and to surface recurring trends, gaps, and design patterns in the literature. To this end, we address the following research questions:
  • RQ1: What are the dominant architectural patterns used to integrate Retrieval-Augmented Generation (RAG) into federated systems?
  • RQ2: What are the primary research focuses and contribution types in the literature on federated RAG systems?
  • RQ3: In which application domains (e.g., healthcare, finance, education) has federated RAG been explored, and what problems are these systems designed to solve?
  • RQ4: What key challenges, open issues, and underexplored areas emerge from current research at the intersection of RAG and federated systems?
These questions structure our analysis of the literature published between 2020 and 2025, allowing us to classify prior work, identify underexplored areas, and provide a foundation for future research at the intersection of privacy, retrieval, and generation.
{forest}
[…]
FedE4RAG (Mao et al., 2025)
FRAG (Zhao, 2024)
, align=left] ] [Dataset / Benchmark, fill=blue!10 [FeB4RAG (Wang et al., 2024b)
RAGRoute (Guerraoui et al., 2025)
, align=left] ] [Evaluation / Analysis, fill=blue!10 [RAGAS (Es et al., 2024), align=left] ] [Survey / Conceptual, fill=blue!10 [FL–LLM surveys (Wu et al., 2025), align=left] ] ] [Application Domain, fill=green!15, tier=intermediate [Healthcare, fill=green!10 [FedRAG (Jiang, 2024)
RAG (Jung et al., 2025)
MIRAGE (Xiong et al., 2024)
, align=left] ] [Finance & Legal, fill=green!10 [C-FedRAG (Addison et al., 2024), align=left] ] [Enterprise / Multilingual, fill=green!10 [MKP-QA (Shojaee et al., 2025), align=left] ] [Recommendation / Personal, fill=green!10 [GPT-FedRec (Zeng et al., 2024), align=left] ] [General / Multi-domain, fill=green!10 [FRAD (Zhao, 2024)
FeB4RAG (Wang et al., 2024b)
, align=left] ] ] [Research Focus, fill=orange!20, tier=intermediate [Privacy & Security, fill=orange!10 [C-FedRAG (Addison et al., 2024)
FRAG (Zhao, 2024)
FedX (Han et al., 2022)
DeTrustFL (Xu et al., 2022)
PPDA (Jeon et al., 2021)
DP-FedKGE (Peng et al., 2021)
, align=left] ] [Retrieval Efficiency, fill=orange!10 [RAGRoute (Guerraoui et al., 2025)
FeB4RAG (Wang et al., 2024b)
, align=left] ] [Model Integration, fill=orange!10 [FedE4RAG (Mao et al., 2025)
UniMS-RAG (Wang et al., 2024a)
FedX (Han et al., 2022)
, align=left] ] [Personalization & Adaptation, fill=orange!10 [GPT-FedRec (Zeng et al., 2024)
MKP-QA (Shojaee et al., 2025)
, align=left] ] ] ]
Figure 1: Classification Scheme: Summarizes the scheme with example categories and the distribution of studies (one study may fall into multiple categories).
2 Methodology
We conducted a systematic literature search to map the research landscape at the intersection of federated systems and retrieval-augmented generation (RAG). Our search spanned top venues in natural language processing (e.g., ACL, EMNLP), machine learning (e.g., NeurIPS, ICLR), security and privacy (e.g., IEEE TIFS), as well as preprint servers like arXiv and relevant industry publications. We used combinations of keywords such as “federated learning”, “retrieval-augmented generation”, and “federated search”, and applied backward reference tracing to ensure inclusion of pivotal works cited in foundational papers (e.g., FedRetriever (Wang et al., 2024a), FLERAG (Mao et al., 2025)).
The search was restricted to works published between 2020 and 2025, reflecting the emergence and evolution of RAG in recent years. Out of an initial set of approximately 50 candidates, we identified 18 primary studies that explicitly addressed the integration of retrieval-based generation into federated or distributed settings. We excluded papers that focused solely on FL or RAG in isolation, without considering their combined architecture, training mechanisms, or deployment context.
Each selected work was systematically categorized along three dimensions, following Kitchenham’s guidelines for systematic mapping studies (Kitchenham et al., 2011): (a) Research Focus: The core technical challenge or system-level objective, such as privacy and security (e.g., secure retrieval, encrypted indexing), retrieval efficiency (e.g., resource selection and latency optimization), model/architecture integration (e.g., coupling RAG with FL pipelines), or domain-specific adaptation (e.g., tailoring systems for healthcare or QA). (b) Contribution Type: The nature of the technical contribution, such as a new model or framework, benchmark dataset, evaluation study, or conceptual survey. (c) Application Domain: The target deployment environment or intended use case, including clinical question answering, legal/financial reasoning, recommender systems, and multi-lingual enterprise settings.
We also recorded the publication year of each study to observe temporal trends and emerging research fronts. Based on this categorization, we developed the unified classification scheme (Figure 1) to summarize the distribution of studies across focus areas and contribution types. Furthermore, this schema informed the design of our architectural taxonomy (Figure 2), which visually organizes the Federated RAG landscape by contribution, application, and system goals.
{forest}
[…]
Architectures, draw, fill=gray!10 [Federated
Retrieval, draw, fill=blue!15 [Naive Aggregation, draw, fill=blue!5] [Selective Query
Routing, draw, fill=blue!5] [Secure Index &
Query, draw, fill=blue!5] ] [Federated
Generation, draw, fill=green!15 [Centralized
Generation, draw, fill=green!5] [Client-specific
Generation, draw, fill=green!5] [Hybrid/Ensemble
Generation, draw, fill=green!5] ] [Privacy/Security
Enhancements, draw, fill=orange!20 [Secure Execution
Environments, draw, fill=orange!10] [Cryptographic
Protocols, draw, fill=orange!10] [Differential Privacy, draw, fill=orange!10] ] ]
Figure 2: Taxonomy of Federated RAG Architectures. A conceptual map to classify the diverse system designs in the literature.
3 Analysis of Recent Studies
Figure 2 categorizes Federated RAG architectures across retrieval strategies, generation mechanisms, and privacy/security enhancements. Retrieval strategies vary from simple naïve aggregation, suitable for small-scale scenarios with trusted participants, to selective query routing (e.g., RAGRoute (Guerraoui et al., 2025)), which reduces redundancy but adds complexity and potential accuracy risks. Secure retrieval using cryptographic protocols (e.g., FRAG (Zhao, 2024)) provides strong privacy protections but introduces computational overhead, making it ideal primarily for highly sensitive contexts.
Generation methods include centralized generation, which ensures consistency but risks privacy breaches and single points of failure, and client-specific generation, which offers superior personalization and data sovereignty at the expense of complexity in maintaining coherence across clients. Hybrid or ensemble strategies attempt to balance these extremes, combining central knowledge with localized adaptations, though increasing management complexity.
These architectural choices require balancing scalability, performance, personalization, and privacy.
3.1 Research Focus
Foundational privacy work from 2020–2022 paved the way for Federated RAG. PPDA (Jeon et al., 2021) introduced lightweight decentralized aggregation without differential privacy (DP) or homomorphic encryption (HE). DP-FedKGE (Peng et al., 2021) added entity-level DP for federated knowledge-graph embedding. DeTrustFL (Xu et al., 2022) used cryptographic consensus to block disaggregation attacks, and FedX (Han et al., 2022) employed cross-client distillation for unsupervised representation learning.
From 2023 through 2025, privacy mechanisms became system-level. C-FedRAG (Addison et al., 2024) executes retrieval and generation entirely inside trusted enclaves, securing clinical QA on MKP-QA. FRAG (Zhao, 2024) realises IND-CPA-secure $k$NN search via single-key HE while keeping latency practical.
Parallel work targets efficiency. RAGRoute (Guerraoui et al., 2025) learns to route queries to high-utility silos, cutting 75 % redundant traffic on MIRAGE. FedB4RAG (Wang et al., 2024b) benchmarks routing policies over 16 BEIR datasets, exposing accuracy–cost trade-offs.
Integration efforts replace isolated modules with end-to-end federated stacks. FedE4RAG (Mao et al., 2025) trains dense retrievers collaboratively, combining HE-protected parameter exchange with distillation to align heterogeneous embeddings. UniMS-RAG unifies client-side retrieval with a shared generator, allowing adaptive control over privacy and latency.
Research Focus # of Studies
Privacy & Security 6
Personalization & Adaptation 2
Retrieval Efficiency 2
Model Integration 3
Table 1: Classification of primary research focus in federated RAG studies, aligned with the taxonomy shown in Figure 1.
Table 1 summarises these trajectories across the 18 primary studies in our map. Half still prioritise privacy and security, reflecting the field’s FL heritage. Personalisation (e.g., GPT-FedRec, MKP-QA) and retrieval efficiency (e.g., RAGRoute, FedB4RAG) are rising, while model-level integration (FedE4RAG, UniMS-RAG) signals a shift toward deployable cross-silo systems.
Overall, research has progressed from algorithmic proofs of concept (2020–2022) to holistic architectures (2023–2025) that simultaneously safeguard data, route queries intelligently, and harmonise retrieval with generation, marking a decisive step toward scalable Federated RAG.
3.2 Contribution Type
The literatures on Federated RAG reflect a balanced evolution across four primary contribution types: new models or frameworks, datasets and benchmarks, evaluation and analysis tools, and conceptual overviews. As shown in Table 2, a substantial portion of the work focuses on developing novel system architectures for privacy-preserving generation and retrieval. FedE4RAG(Mao et al., 2025) introduces a modular design that integrates homomorphic encryption and knowledge distillation to train federated dense retrievers, offering a decentralized alternative to traditional centralized systems. C-FedRAG(Addison et al., 2024) builds a confidential execution pipeline using Trusted Execution Environments (TEEs), enabling secure generation and retrieval particularly suited to sensitive domains like healthcare. FRAG (Zhao, 2024) contributes a cryptographic kNN retrieval system backed by IND-CPA-secure homomorphic encryption, facilitating privacy-preserving vector search.
Beyond system-level proposals, several contributions focus on datasets and benchmarking. FeB4RAG(Wang et al., 2024b) presents the first federated retrieval benchmark across 16 BEIR-derived domains, enabling rigorous evaluation of routing strategies and retrieval quality under resource constraints. MKP-QA(Shojaee et al., 2025) introduces a domain-specific benchmark for enterprise multilingual QA, designed to assess cross-silo document selection and context preservation. Additionally, RAGRoute (Guerraoui et al., 2025) offers a specialized evaluation testbed for its adaptive query planning mechanism, allowing performance measurement across silos under constrained budgets.
Contribution Type # of Studies
New Model / Framework 3
Dataset / Benchmark 2
Evaluation / Analysis 1
Survey / Conceptual 1
Table 2: Categorization of studies by primary contribution type. Majority propose new models; others contribute benchmarks, evaluation metrics, or conceptual overviews.
Evaluation and analysis contributions aim to establish consistent standards for measuring retrieval and generation quality in federated setups. RAGAS (Es et al., 2024) proposes a suite of evaluation metrics tailored to federated RAG pipelines, accounting for inconsistencies in retrieved evidence, hallucinations, and robustness to noise injection. FeB4RAG extends this line of work by benchmarking retrieval degradation when shifting from naive to intelligent routing policies, illustrating the tradeoffs between accuracy and resource consumption.
Finally, survey and conceptual literature remains sparse. The only entry in this category is FL-LLM Surveys (Wu et al., 2025), which provides a broader discussion on federated LLMs and briefly references RAG mechanisms. In contrast, our mapping study represents to the best of our knowledge, the first structured synthesis of architectural, empirical, and deployment-oriented work in Federated RAG, addressing a clear gap in the literature.
3.3 Application Domains and Use Cases
Federated RAG is gaining traction across a variety of domains that require both strong privacy guarantees and access to distributed knowledge sources. Among these, healthcare has emerged as the most prominently explored application vertical, driven by high data sensitivity and institutional silos. Jung et al. (2025) demonstrated that federated RAG pipelines outperform traditional FL setups on clinical question answering tasks by retrieving from a shared medical literature corpus while preserving patient privacy. Complementing this, Jiang (2024) introduced a hierarchical retrieval approach that selects relevant hospitals before accessing localized documents, facilitating fine-grained medical inference. MIRAGE (Xiong et al., 2024) further formalizes this direction through a clinical QA benchmark tailored for federated evaluation, underscoring the domain’s methodological maturity.
Outside of healthcare, finance and legal sectors have been identified as strong candidates for Federated RAG due to similar compliance and confidentiality constraints. C-FedRAG(Addison et al., 2024) targets regulation-aware summarization and document generation, although large-scale evaluation in these domains remains limited. In enterprise and multilingual QA, MKP-QA(Shojaee et al., 2025) introduces a federated benchmark for retrieving internal product knowledge across silos and languages without centralizing sensitive documentation. Meanwhile, GPT-FedRec (Zeng et al., 2024) explores personalized recommendation using a hybrid framework that combines collaborative filtering with retrieval-augmented generation, showing reductions in hallucination and data sparsity.
Application Domain # of Studies
Healthcare 3
Finance & Legal 1
Enterprise / Multilingual 1
Recommendation / Personal 1
General / Multi-domain 2
Table 3: Application domains addressed in federated RAG research. Studies are counted by explicit domain targeting, with several works remaining domain-agnostic or generalized.
In addition to these domain-specific efforts, several studies adopt general or multi-domain settings. FeB4RAG(Wang et al., 2024b) provides a comprehensive evaluation across 16 BEIR-derived tasks, while FRAG(Zhao, 2024) focuses on architectural generalization under cryptographic constraints. These studies emphasize benchmarking and scalable deployment over specific industry contexts.
As summarized in Table 3, healthcare dominates the literature with three focused studies, followed by singular works in finance/legal, enterprise/multilingual QA, and recommendation systems. This distribution aligns with real-world concerns over privacy, compliance, and heterogeneous data silos. Earlier Figure 2 illustrated how such deployment settings are reflected in the system-level architecture choices across the literature. While many models are still domain-agnostic, healthcare and enterprise NLP tasks are emerging as leading targets for practical Federated RAG adoption, with future work likely to expand into education, public sector governance, and legal tech.
4 Research Activity and Publication Trends
Research in Federated Retrieval-Augmented Generation (RAG) has accelerated notably in recent years, driven by rising demand for privacy-preserving NLP and the maturation of retrieval-augmented systems. Before 2022, federated learning (FL) and RAG evolved independently, FL focused on decentralized optimization and privacy, while RAG addressed hallucinations in large language models (LLMs) (Lewis et al., 2020a). Their integration began in 2022–2023 with early architectural proposals such as UniMS-RAG (Wang et al., 2024a) and FedE4RAG (Mao et al., 2025), which outlined how to combine distributed training with retrieval-based augmentation.
As shown in Figure 3, publication volume increased significantly in 2024, with at least eight new studies emerging across venues in NLP (e.g., Findings of ACL), ML systems (e.g., MLSys, NeurIPS workshops), secure computing (e.g., IEEE TIFS), and industry white papers. This surge is not just numerical but also reflects broader shifts discussed in previous sections, specifically the diversification of research focus (Table 1) and emergence of novel models, benchmarks, and real-world applications (Table 2, Table 3). The steep rise in 2024 corresponds to two converging trends:
(i) The expanding use of LLMs in high-stakes domains such as healthcare and finance, where hallucination and outdated information are critical concerns.
(ii) Increasing regulatory and ethical scrutiny over centralized data processing, which incentivizes federated architectures.
Despite this momentum, the limited number of relevant studies observed in 2023 likely reflects the early-stage nature of Federated RAG research at that time, characterized by initial exploration of architectural concepts and limited practical deployments, preceding the subsequent significant increase in publication activity observed from 2024 onward, mostly appeared as preprints or workshop papers, with limited presence in archival venues. Notably, MIRAGE (Xiong et al., 2024) became one of the first benchmarks accepted to Findings of ACL 2024, signaling early consolidation around evaluation protocols. This aligns with trends seen in earlier paragraph on contribution type 3.2, where evaluation and analysis are gaining traction alongside system development.
Industry engagement is also increasing. C-FedRAG  (Addison et al., 2024) from NVIDIA proposes TEE-based confidential QA, while Adobe’s MKP-QA  (Shojaee et al., 2025) focuses on multilingual enterprise retrieval. Meanwhile, open-source initiatives like the Vector Institute’s FedRAG toolkit point to growing developer interest in practical deployment. Together, these developments reinforce the trends visualized in Figure 3, validating the observations made in the application domain and contribution type analyses. We expect Federated RAG to evolve into a distinct research track within NLP and AI by 2025–2026, underscoring the relevance of this systematic mapping study for guiding future work.
20202021202220232024202500555510101010# of Publications
Figure 3: Publication trends on federated RAG from 2020 to 2025, showing a significant rise in research activity in recent years.
4.1 Privacy & Security
Federated RAG must secure both retrieval and generation. C-FedRAG confines the entire pipeline to SGX enclaves, shielding queries and documents from untrusted servers. FRAG encrypts vector search with IND-CPA–secure homomorphic $k$NN, avoiding raw-index exposure, while FedE4RAG performs local augmentation at inference to minimise data movement. These techniques uphold regulations such as “right-to-be-forgotten” but add latency and can hamper cross-silo reasoning; future work should formalise threat models and quantify the privacy–utility frontier.
4.2 Model Adaptation & Knowledge Management
Maintaining knowledge without a central store is equally demanding. FedE4RAG trains dense retrievers federatively, preserving locality, while adapter-style updates and federated embedding learning cut communication. Hybrid planners like RAGRoute select local or remote answers when indices drift. Yet index synchronisation, conflict resolution, and continual adaptation across heterogeneous silos remain open. CRDT-based distributed indexing and meta-learned retriever personalisation are promising, but benchmarks must cover timestamped document shifts, bandwidth ceilings, and real-world constraints.
5 Applications
Federated RAG is particularly suited for domains where data sensitivity, personalization, and access to localized knowledge are paramount. In healthcare, it enables clinical question answering systems that draw on hospital-specific knowledge bases while ensuring that sensitive patient data remains on-premise, complying with strict privacy regulations. In financial services, it facilitates regulation-aware summarization and reporting by incorporating institution-specific documents and compliance guidelines without exposing confidential data externally. Education, enterprise knowledge management, and legal technology also present high-impact opportunities. In education, federated RAG can power personalized tutoring systems that align with local curricula and student learning profiles. Enterprise use cases include internal document retrieval and report generation without requiring the centralization of proprietary content. Legal applications range from case-specific document analysis to retrieval-augmented reasoning over firm-confidential corpora.
These examples illustrate the broad utility of federated RAG in supporting privacy-preserving, context-aware, and domain-specific language generation across distributed settings.
6 Challenges, Gaps and Trends
In this section, we discuss the overarching challenges, research gaps, and emerging trends for federated RAG. Table 4 summarizes how our research questions (RQ1–RQ4) align with specific challenge areas identified from the 18 primary studies.
Research Question Key Insights / Challenge Areas
RQ1 System heterogeneity in data and clients; need for scalable federated RAG architectures.
RQ2 Privacy and security trade-offs; mechanisms for trust and compliance in federated setups.
RQ3 Retrieval and generation efficiency; balancing quality against resource costs.
RQ4 Evaluation gaps and future directions, such as benchmarks and domain-specific adaptations.
Table 4: Mapping of RQ1–RQ4 to key insights and challenge areas in federated RAG.
System Heterogeneity, Scalability, and Performance Trade-offs – Federated RAG systems must navigate heterogeneity in data distribution, storage capacity, and computational power across clients. Unified retrieval pipelines face challenges due to non-IID data and inconsistent resource availability. Naively querying all silos introduces latency and noise. To address this, RAGRoute employs lightweight classifiers for selective query routing, reducing source queries by up to 75% while maintaining answer quality. Similarly, the FeB4RAG benchmark highlights the importance of intelligent source selection and merging strategies in improving retrieval efficacy. Managing performance trade-offs, such as recall versus latency, is critical. Recent work like FRAG introduces encrypted caching and probabilistic gating to optimize resource use while maintaining privacy. Overall, these advances address architectural and efficiency concerns central to scalable federated RAG (RQ1).
Privacy and Security Considerations – Federated RAG must guarantee data confidentiality throughout both retrieval and generation, making privacy a core concern (RQ2). Traditional RAG pipelines risk information leakage through exposed queries or retrieved documents. Recent systems address this through trusted execution and encryption. C-FedRAG, for example, runs indexing and generation entirely inside secure enclaves, protecting sensitive content even when orchestrated by untrusted servers. FRAG further strengthens guarantees via homomorphic encryption, enabling encrypted vector search without exposing embeddings or indices—achieving IND-CPA security with acceptable overhead. Compliance with data governance (e.g., right-to-be-forgotten) motivates designs that avoid raw data centralization. FedE4RAG exemplifies this by performing local augmentation at inference, ensuring data minimization. However, privacy-preserving techniques often introduce latency and limit cross-silo reasoning, raising the challenge of balancing security with system utility. Future work must formalize threat models and quantify trade-offs between confidentiality, accuracy, and responsiveness in real-world deployments.
Architectural Maturity – Retrieving and updating knowledge without centralized storage is a central challenge for federated RAG systems (RQ3). FedE4RAG proposes client-side retriever training via federated learning, improving generalization while respecting data locality. To mitigate resource constraints, current approaches favor parameter-efficient methods such as adapter tuning or federated embedding learning. Hybrid pipelines like RAGRoute arbitrate between federated and RAG responses based on confidence, improving robustness when document indices are stale. However, synchronizing indices, resolving data conflicts, and enabling continual adaptation across diverse silos remain open problems. For domain-specific systems, e.g., hospital QA linked to EHR silos, federated RAG must reconcile personalization with evolving knowledge, potentially through distributed indexing, meta-learning, and lightweight model updates. Future work should explore CRDT-based index synchronization and meta-learned retriever personalization to enable continual, privacy-preserving adaptation across heterogeneous silos. Benchmarking these techniques under realistic, timestamped document shifts and resource constraints remains an open evaluation priority.
Evaluation and Benchmarking – Federated RAG is a new topic, hence evaluation frameworks (RQ4) are lacking. Early studies reused QA criteria or imitated federated environments, making comparisons impossible. Federated retrieval and creation require standardized datasets and metrics to account for client relevance, privacy constraints, and network costs. This gap is being filled by recent work: FeB4RAG provides a dataset for federated RAG search, allowing for realistic multi-silo evaluation of retrieval algorithms. FedRAG may leverage the MIRAGE benchmark (used by RAGRoute) and a new MIMIC-IV-based clinical QA dataset to evaluate systems in certain areas. Still, evaluating issues exist. Relevance across clients requires reinterpreting IR measurements like recall or MRR. Current measurements make it difficult to measure a system’s privacy-preserving quality (utility attained for a given level of privacy). Real-world deployment studies are rare; most are lab-based. Common evaluation protocols, such as federated versions of popular QA tasks and defined success criteria (accuracy vs. bandwidth vs. privacy), will advance the industry. Community-driven toolkits (e.g., Flower FedRAG demonstrations) and open-source frameworks are lowering experimental barriers, which are beneficial. Future work should introduce privacy–utility Pareto benchmarks that vary attacker models and bandwidth ceilings, alongside live-stream leaderboards that track adaptation latency and cost as documents drift over time.
6.1 Outlook and Future Opportunities
Federated RAG research is rapidly evolving, addressing challenges in privacy, scalability, and personalization. Key opportunities include reducing cross-silo latency via neural routing and CRDT-based index synchronization, strengthening privacy through Pareto benchmarks that jointly track attacker models and bandwidth ceilings, and enabling continual adaptation with meta-learned retriever personalization and tiny-ML adapters scheduled to respect heterogeneous devices. Lifelong evaluation resources such as live-stream leaderboards and federated versions of popular QA tasks will clarify trade-offs among accuracy, privacy, and cost as documents drift. Domain-specific optimizations for healthcare, finance, and multi-enterprise settings remain crucial, and community toolkits are lowering experimental barriers. Overall, federated RAG offers a compelling path toward trustworthy, knowledge-grounded AI that operates securely in decentralized environments. Future research could empirically investigate and quantify these identified architectural trade-offs across diverse real-world federated contexts, thereby enabling more precise architectural recommendations tailored explicitly to domain-specific operational constraints.
7 Conclusion
Federated Retrieval-Augmented Generation (RAG) has rapidly moved from concept to practice by pairing privacy-preserving training with knowledge-grounded generation. Our systematic mapping of 18 primary studies charts this evolution, classifying architectures, contributions, and domains, and highlighting a 2024 research surge driven by large-scale language-model adoption and regulatory pressure for confidential data handling.
Where it already helps – Early deployments demonstrate clear value in compliance-sensitive settings: C-FedRAG answers clinical questions while keeping electronic health-record data in-hospital; FedE4RAG powers enterprise knowledge assistants without centralizing proprietary documents; and RAGRoute cuts 75% of redundant cross-silo queries, lowering latency for finance workloads. Across these use cases, federated RAG improves factual grounding, reduces hallucinations, and upholds data-sovereignty requirements that vanilla RAG or centralized LLM serving cannot meet.
What has been solved and what has not – Secure enclaves, homomorphic encryption, and differential-privacy mechanisms now offer end-to-end confidentiality; neural routing and probabilistic caching tame bandwidth costs; and federated embedding learning supplies parameter-efficient personalization. Yet four hurdles persist:
  • Scalable, CRDT-style index synchronization to keep silos consistent,
  • Meta-learned retriever adaptation that survives concept drift,
  • Privacy–utility evaluation protocols that expose trade-offs under realistic attacker models, and
  • Live benchmarks that measure cost and latency as documents evolve.
Why this mapping study matters – Our study comprehensively maps the developing area of federated RAG to identify the technical gaps that hinder application in healthcare, banking, education, and cross-enterprise cooperation, where data sovereignty is non-negotiable but knowledge sharing is crucial. These problems require shared benchmarking testbeds, robust threat models, and lightweight on-device adapters to expand access to federated RAG to enable trustworthy AI in these environments. Federated RAG can grow from a promising niche to a foundation for secure, knowledge-rich language technology by bridging NLP, security, and distributed systems.
Limitations
As with any systematic mapping study, our work is subject to certain limitations that should be acknowledged when interpreting its findings.
External Validity
This study aimed to synthesize trends and insights in Federated RAG research based on a carefully defined corpus of 18 primary studies spanning 2020–2025. To ensure external validity, we explicitly scoped our review to works that integrate retrieval-augmented generation within federated or privacy-preserving systems, across NLP, IR, ML, and security venues. While we employed backward reference tracing and cross-domain search strategies, there remains a possibility that relevant studies, especially those in proprietary industry settings or less commonly indexed venues, may have been missed. Furthermore, the current corpus reflects research trends primarily from Western and East Asian academic and industrial ecosystems, and may not fully represent globally distributed research efforts.
Construct Validity
Federated RAG is an emerging area, and its boundaries are still fluid across communities. Although we adhered to a strict inclusion protocol and used Kitchenham’s guidelines for mapping studies, differing interpretations of what constitutes a Federated RAG system (e.g., systems using static knowledge bases vs. dynamic retrieval) may introduce conceptual variance. In some cases, papers labeled as federated learning systems implicitly used retrieval or augmentation mechanisms without explicit reference to RAG paradigms. While we erred on the side of inclusivity, such conceptual ambiguity may influence the completeness or framing of our categories. Moreover, reliance on peer-reviewed and preprint literature means that grey literature, technical documentation, or non-archival systems (e.g., deployed prototypes) were excluded.
Conclusion Validity
Our findings are based on a targeted but relatively small number of primary studies. While this reflects the nascent state of the field, it also means that statistical generalizations or trend extrapolations should be interpreted with caution. For instance, the observed increase in 2024 publications may reflect a short-term spike due to LLM popularity rather than a long-term trajectory. Additionally, the evolving definitions and tooling of federated learning and RAG (e.g., hybrid decentralization models, model-serving architectures) suggest that future iterations of the field may adopt different terminology or structural assumptions. As such, the taxonomy and classifications presented here represent a snapshot of a fast-changing landscape.
Other Limitations
Selection bias remains a potential risk: our search protocol, while comprehensive, may have inadvertently excluded studies that use unconventional terminology or fall at the periphery of federated and retrieval-based learning. Furthermore, because many contributions in this domain come from industry, the absence of open-source implementations or full technical disclosures limits reproducibility and may skew the perceived landscape toward academic publications. Finally, while we have captured the most salient developments up to early 2025, the pace of innovation in LLM-based systems warrants periodic re-evaluation to maintain relevance and coverage.
Acknowledgment
We gratefully acknowledge the Complex Data Analysis and Reasoning Lab at Arizona State University for their resources and computational support.